evaluate sin43 degree using Taylor's expansion
Answer by Alex.33(110)

Taylor's series: Given x is close to a,
Since you are estimating, take the first three above, and set a=pi/4, so x-a=-2 deg=-pi/90.
=0.6819933...
Approximately it will be happy to be 0.682.
BTW the difference between this approximation and the exact value is contained within 5.06*10^-6.
And there're other fast ways to do that.
>Apply Maclaurin's series instead(take a=0). A lot faster.
>A much more simple method: sin(43deg)=sin(pi/4)-(sin(pi/4))'*(pi/90). Super fast and you'll be impressed by its accuracy.